# Spin polarized enantio-sensitive multipolar photoelectron currents

### v3

@@ Abstract
Photoelectron circular dichroism (PECD) manifests as a forward-backward asymmetry of electron emission in the direction orthogonal to the light polarization plane via one-photon ionization of chiral molecules with circularly polarized light. Multi-polar ‘PECD’ currents, i.e., currents resolved along multiple directions, have also been predicted using two mutually-orthogonal linearly polarized light with carrier frequencies $\omega$ and $2\omega$. These currents arise from the interference between the one- and two-photon transitions. Here, we will show that photoelectron spin detection already reveals enantio-sensitive multi-polar currents in the one-photon regime since the two axes can be marked by the photoelectron momentum $\boldsymbol{\hat{k}}$ and spin-detection axis $\boldsymbol{\hat{s}}$. Specifically, we consider one-photon ionization of an isotropic ensemble of randomly oriented chiral molecules and show that the direction of the resulting photoelectron current is enantio-sensitively ‘locked’ to the photoelectron’s spin, which is mediated by two mechanisms. First, is the Bloch pseudovector which enables a collinear locking forming either a spin-sink or source for opposite enantiomers. Second, is the spin torque pseudovector that enables orthogonal locking forming a spin vortex in the polarization plane that rotates in opposite directions for opposite enantiomers. The former effect is a spin analog of photoelectron vortex dichroism (Phys. Rev. Lett. 129, 233201, 2022) wherein the detected photoelectron spin encodes molecular chirality while the latter is reminiscent of the Rashba effect in solids.
Photoelectron circular dichroism (PECD) heralded the “dipole revolution” in chiral discrimination: chiral discrimination without using chiral light [25, 5]. It was first predicted by Ritchie [31] and first detected by Böwering et. al. [6]. Nowadays, PECD is very well established theoretically, and also shown to yield strong enantiosensitive signals across several molecular species [29, 17, 37, 18, 21, 16, 19, 20, 22, 35, 11, 30, 36, 4]. In PECD, the photoionization of an isotropic ensemble of randomly oriented chiral molecules via circularly polarized light results to a forward-backward asymmetry in the net photoelectron current in the direction of light polarization. Fundamentally, it can be understood as a manifestation of a geometric magnetic field introduced in Ref. [25], and its emergence leads to new enantio-sensitive observables in photoionization [5, 23].
Geometric magnetism in photoionization [23] and photoexcitation [27, 32] of chiral molecules addresses the dynamical origin of enantiosensitive observables in one or multiphoton ionization. Its central object is the geometric propensity field:
$\boldsymbol{\vec{B}}_{\boldsymbol{\vec{k}}}\equiv i\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}}^{*}\times\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}},$ (1)
where, $\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}}}$ is the photoelectron dipole field. The propensity field underlies several classes of such observables that originate from its multipole moments [23]: (i) the net propensity field $\boldsymbol{\vec{\Omega}}\equiv\int d\Theta_{k}\boldsymbol{\vec{B}}_{\boldsymbol{\vec{k}}}$; (ii) the net radial component of the propensity field $B^{\parallel}\equiv\int d\Theta_{k}(\boldsymbol{\hat{k}}\cdot\boldsymbol{\vec{B}}_{\boldsymbol{\vec{k}}})$; and (iii) the spherical multipole moments of the longitudinal $B_{l,m}^{\parallel}\equiv\int d\Theta_{k}(\boldsymbol{\hat{k}}\cdot\boldsymbol{\vec{B}}_{\boldsymbol{\vec{k}}})Y_{l,m}$ and transversal field components , $B_{l,m}^{\perp,1}\equiv\int d\Theta_{k}\boldsymbol{\vec{B}}_{\boldsymbol{\vec{k}}}\cdot\nabla_{k}Y_{l,m}$, and $B_{l,m}^{\perp,2}\equiv\int d\Theta_{k}\boldsymbol{\vec{B}}_{\boldsymbol{\vec{k}}}\cdot(\boldsymbol{\hat{k}}\times\nabla_{k})Y_{l,m}$.
PECD belongs to Class II observables [24], .i.e, $\boldsymbol{\vec{j}}_{\text{PECD}}\propto B^{\parallel}$, while Class I observables have been explored in Ref. [23]. It was shown that the net propensity field $\boldsymbol{\vec{\Omega}}$ controls the enantio-sensitive orientation of the cations - an effect called molecular orientation circular dichroism (MOCD). Class III observables have tensorial nature and represent multipolar photoelectron currents, i.e., currents resolved along multiple directions. The properties of such observables emerging in two photon ionization by two-color fields due to the interference between the one- and two-photon transitions have been explored theoretically in Refs. [13, 14, 26]. The same two-color set-up enabled detection of its possible analogue in multiphoton regime [33]. However, these effects ignore the photoelectron spin.
Spin detection opens an exciting opportunity to reveal enantio-sensitive multipolar currents arising already in the one-photon regime. Indeed, the two axes needed to resolve such currents can be marked by the photoelectron momentum $\boldsymbol{\hat{k}}$ and spin-detection axis $\boldsymbol{\hat{s}}$, see Fig. 1. The analysis of the interplay of chirality and spin in photoionization of chiral molecules was pioneered by Cherepkov [9, 10, 12]. He identified a kinematic picture of spin polarization, predicting spin- and enantio-sensitive effects in one photon ionization for circularly and linearly polarized light. First experiments detecting spin polarization in multiphoton ionization have just appeared [2]; simulations predicting spin-polarization in PECD have also become possible [3].
@@
Here, we use our formalism on spin-resolved one-photon photoionization of chiral molecules [15] to explore multipolar photoelectron currents, and show that these currents are enabled by two complementary geometric mechanisms that arise solely from electric dipole interactions and the geometric properties of the photoionization dipoles in real space and spin space. The first mechanism is mediated by the Bloch pseudovector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$, and operates in fields of arbitrary polarization, including fully isotropic polarization. The second mechanism is the spin-resolved propensity field $\vec{\mathbb{B}}_{\boldsymbol{\vec{k}}}$ which is a natural extension of Eq. (1) and is now a matrix in spin space. It only operates in fields with polarization planes fixed in the laboratory frame, e.g., circularly or elliptically polarized light.
@@
We also use the same synthetic chiral argon system employed in Ref. [15], which are constructed by combining excited-state orbitals:
@@
These sates $|\psi_{m,\mu}^{\pm}\rangle_{p}$ and $|\psi_{m,\mu}^{\pm}\rangle_{c}$ are inspired by analogous chiral hydrogenic states [25]. Unlike hydrogen, the multielectron core potential in argon breaks inversion symmetry but the synthetic chirality is stabilized by electron correlations. The resulting photoelectron currents are equivalent to earlier predictions of Cherepkov [9, 10, 12], allowing us to identify the dynamical origins of spin polarization in photoionization. The direct comparison of our approaches will be addressed elsewhere.
@@
The photoelectron current conditioned on the spin detection axis $\boldsymbol{\hat{s}}^{L}$ can now be calculated as follows:
@@
wherein, $\int d\rho$, $\int d\Theta_{k}^{L}$, and $\int d\Theta_{s}^{L}$ denotes averaging over all molecular orientations, photoelectron momentum, and spin detection axis, respectively. Full details of the succeeding calculations are shown in Appendix A.
Let us first consider randomly oriented chiral molecules under isotropic illumination by linearly polarized light, i.e., instead of having a single direction $\boldsymbol{\hat{\epsilon}}^{L}$ to characterize to polarization direction (see Fig. 1), we introduce
$\boldsymbol{\vec{E}}^{L}=$ $E_{\omega}^{L}\left(\sin\theta_{p}\cos\varphi_{p}\boldsymbol{\hat{x}}^{L}+\sin\theta_{p}\sin\varphi_{p}\boldsymbol{\hat{y}}^{L}+\cos\theta_{p}\boldsymbol{\hat{z}}^{L}\right).$ (8)
Subsituting Eq. (8) into (7) and averaging over all polarization directions $\int d\Theta_{p}$, the resulting photoelectron current is now
@@
where, $S_{0}$ is the total yield
$S_{0}=\int d\Theta_{k}\left(\left|\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},\frac{1}{2}}\right|^{2}+\left|\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},-\frac{1}{2}}\right|^{2}\right).$ (10)
Equation (9) shows that the current is proportional to the flux of the vector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}$ through the surface of the energy shell ($d\boldsymbol{\vec{\Theta}}_{k}^{M}=d\Theta_{k}^{M}\boldsymbol{\hat{k}}^{M}k^{2}$). Mathematically, $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ is a Bloch pseudovector:
@@
describing spin orientation in the degenerate two-level system formed by spin-up and spin-down continuum states populated by photoionization, which incorporates both populations and coherences within this system [15]. Here, $\tilde{\varrho}$ is the reduced density matrix which emerges after performing a partial trace over the spatial continuum states and degenerate ionization channels of the full density matrix, and averaging over all molecular orientations.
@@
and that the current is collinearly ‘locked’ to the direction of the spin-detection axis $\boldsymbol{\hat{s}}$. Hence, $\boldsymbol{\vec{j}}_{\text{iso}}^{L(S)}(\boldsymbol{\hat{s}}^{L})$ serves as the spin analog of photoelectron vortex dichroism (PEVD) [28]. Since, PEVD uses linearly polarized light to generate photoelectrons with orbital angular momentum (OAM) wherein the OAM has opposite signs for opposite enantiomers, Eq. (9) now suggests that this collinearly ‘locked’ photoelectron current can be enhanced by using light fields with either a well-defined polarization direction or polarization plane.
Figure 2: Schematic of the resulting photoelectron currents for (a,b) linearly, and (c-f) circularly polarized light in the laboratory frame. The solid and dashed blue vectors denote the opposite directions of spin for opposite enantiomers. (a-d) The currents $\boldsymbol{\vec{j}}_{\text{lin}}$ and $\boldsymbol{\vec{j}}_{r}$ are enabled by the Bloch vector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ which results to a collinear locking of the photoelectron current and spin $\boldsymbol{\vec{j}}\parallel\boldsymbol{\hat{s}}$. (e-f) The currents $\boldsymbol{\vec{j}}_{\text{PECD}}$ and $\boldsymbol{\vec{j}}_{\tau}$ are enabled by the spin-resolved propensity field $\vec{\mathbb{B}}_{\boldsymbol{\vec{k}}}$. The PECD current is not spin-sensitive, while $\boldsymbol{\vec{j}}_{\tau}$ presents a three-way orthogonal locking of the photoelectron current and spin with the photon spin.
Repeating the same process for linearly polarized light
@@
The additional vector $\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}$ is the directional bias introduced by the well-defined direction of the light polarization. Hence, the photoelectron current now arises as the flux of an effective Bloch pseudovector $\boldsymbol{\vec{\mathsf{S}}}=a\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}+b\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}$ through the surface of the energy shell.
Figure 3: Comparison of the photoelectron currents enabled by the Bloch vector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ for different chiral states [Eqs. (2)-(3)] via random illumination, linearly (a,b) and circularly polarized light (c,d) with either orthogonal or collinear detection geometry [see Fig. 1]. The rapidly oscillating behavior at higher values of $k$ are due to the Fano resonances, leading up to the ionization threshold for the 3s electrons [34, 8]
The current $\boldsymbol{\vec{j}}_{\text{lin}}^{L}$, Eq. (14), can be measured using two possible schemes: (i) collinear $\boldsymbol{\hat{s}}\parallel\boldsymbol{\hat{\epsilon}}$ and (ii) orthogonal $\boldsymbol{\hat{s}}\perp\boldsymbol{\hat{\epsilon}}$ detection geometry, see Fig. 2(a,b), respectively. Enhancement or reduction of $\boldsymbol{\vec{j}}_{\text{lin}}^{L}$ with respect to $\boldsymbol{\vec{j}}^{L}_{\text{iso}}$ is then controlled by $\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}$:
@@
A comparison of the currents $\boldsymbol{\vec{j}}_{\text{iso}}^{L}$ and $\boldsymbol{\vec{j}}_{\text{lin}}^{L}$ for the states $|\psi_{+1,\frac{1}{2}}^{\pm}\rangle_{c}$ and $|\psi_{-1,\frac{1}{2}}^{\pm}\rangle_{p}$ is shown in Fig. 3(a,b). It can be seen that enhancement of the current is indeed possible but not always ensured as it is sensitive both to the chiral state and detection geometry. Figure 3(a) demonstrates that the orthogonal detection geometry enhances the photoelectron current but is reduced for collinear detection geometry. Meanwhile, Fig. 3(b) shows that both detection geometries enhances the resulting photoelectron current with the collinear detection geometry now having a larger enhancement.
@@
The term $\boldsymbol{\vec{j}}_{r}^{L}$ is enabled by the Bloch vector $\boldsymbol{\vec{\mathcal{S}}}^{M}_{\boldsymbol{\vec{k}}^{M}}$ and has a similar form as that of Eq. (14):
@@
which is enantio-sensitive but not dichroic. The collinear locking between the photoelectron current and spin is still present can be measured either via collinear or orthogonal detection geometry with respect to the photon spin direction $\boldsymbol{\hat{\Xi}}^{L}=(-i\boldsymbol{\vec{E}}^{L*}\times\boldsymbol{\vec{E}}^{L})/|\boldsymbol{\vec{E}}^{L}|^{2}=\xi\boldsymbol{\hat{z}}^{L}$, see Fig. 2(c,d). Moreover, the current $\boldsymbol{\vec{j}}_{r}^{L}$ is contained in the coefficients $B_{1}$ and $B_{2}$ introduced by Cherepkov [9, 10, 12]. Figure 3(c) shows that there is negligible effect to the current compared to $\boldsymbol{\vec{j}}_{\text{iso}}^{L}$. Meanwhile, Fig. 3(d) demonstrates enhancement of the current but it is almost half of $\boldsymbol{\vec{j}}_{\text{lin}}^{L}$.
@@
which is now a matrix in spin space, and a natural extension of Eq. (1). Explicitly, we have
@@
which shows that PECD is not spin sensitive since it takes the sum of the spin-up and down photoelectrons, and proportional to the flux of the propensity field through the surface of the energy shell ($d\boldsymbol{\vec{\Theta}}^{M}=d\Theta_{k}^{M}\boldsymbol{\hat{k}}^{M}k^{2}$). Equation (20) is equal to the PECD current we derived in [24], to the original expression derived by Ritchie [31], and to the coefficient $D$ introduced by Cherepkov [9, 10, 12].
Figure 4: (a,b) Comparison of the components of $\boldsymbol{\vec{j}}_{\text{circ}}$ for different chiral states [Eqs. (2)-(3)] with $\boldsymbol{\hat{s}}\parallel\boldsymbol{\hat{y}}$.
@@
is confined in the polarization plane which indicates transversal spin-polarization. This current is maximal when the photoelectron spin detection axis $\boldsymbol{\hat{s}}^{L}$ is orthogonal to both the direction of the current and photon spin $\boldsymbol{\hat{\Xi}}^{L}$. Equation (21) can also be equivalently written as the flux of the spin torque vector $\boldsymbol{\vec{\tau}}_{\boldsymbol{\vec{k}}^{M}}^{M}$ through the surface of the energy shell. It is also equivalent to the coefficient $C$ of Cherepkov [9, 10, 12], in which $\boldsymbol{\vec{j}}_{\tau}^{L}\propto A^{1,1}_{1,-1}-A^{1,-1}_{1,1}$. The enantio-sensitive nature of $\boldsymbol{\vec{j}}_{\tau}^{L}$ arises from the molecular factor $(\boldsymbol{\vec{\tau}}^{M}\cdot\boldsymbol{\vec{k}}^{M})$ which is a pseudoscalar: opposite enantiomers produce opposite spin vortices, such that spin polarization direction depends on photoelectron momentum in a way reminiscent of Rashba effect in solids.

### v4

# arXiv:2505.23460v4[quant-ph] 26 Feb 2026

@@ Abstract
Chiral molecular structures enable time-odd pseudotensor that controls correlations between photoelectron spin and molecular orientation in photoionization [13]. We show that chiral structures also support time-even correlations between the photoelectron spin and its momentum. Fundamentally, these correlations can only be revealed in conditioned measurements. We argue that conditioned measurements are the sole origin of a broad class of phenomena termed chirality-induced spin selectivity (CISS), because all these phenomena require a correlated detection of the spin and another variable, e.g., electron current. We consider one-photon ionization of an isotropic ensemble of randomly oriented chiral molecules and show that the direction of the resulting photoelectron current is enantio-sensitively ‘locked’ to the photoelectron spin even under isotropic illumination. We also identify molecular frame spin texture created by photoionization in which its flux through the energy shell is the molecular pseudoscalar that quantifies the underlying spin-current correlations. Next, we consider the effect of the photon spin on these correlations and find that chiral structure also leads to triple correlations between the photoelectron momentum, its spin and the photon spin. Our results identify two molecular pseudovectors underlying the full complexity of the multitude of spin-conditioned measurements of the photoelectron current
Chirality induced spin selectivity (CISS) usually refers to a spin-polarized charge transport through chiral molecules [29], but in a wider context it underscores a recently discovered functionality of chiral molecules to filter photoelectron spin [12]. CISS is surrounded with intrigue and debates on its origin and strength which stimulate detailed simulations, simple models and new experiments [12]. CISS can be initiated either by photoionization [16] or by applying a voltage across a chiral molecular device [36]. The latter usually involves a substrate coated with a thin film of chiral molecules and a spin-sensitive probe, such as a magnetic conductive probe atomic force microscope (mCP-AFM). The tip of the microscope is magnetized to resolve the current conditioned on spin. Flipping the direction of tip’s magnetisation, one detects the difference in the current through the device.
This example shows that CISS related phenomena are fundamentally linked to conditioned measurements. Such measurements reveal correlations between some vectorial molecular property $\mathbf{v}$ and an electron spin $\mathbf{s}$: $\langle\mathbf{v}\otimes\mathbf{s}\rangle$. Such pseudoscalar correlations can only be enabled by chiral structures and can have different strength and manifestations depending on the specific measurement. Starting at the premises of conditioned measurements, we consider the simplest possible scenario of chiral spin-filtering: photoionization of an ensemble of randomly oriented chiral molecules. Our aim is to identify the origin of such pseudoscalar correlations. The simple scenario that has not yet been quantified is the one removing all external biases such as molecular orientation or polarization of the light field.
Consider one photon ionization of randomly oriented chiral molecules under isotropic illumination. To detect a net current, one needs to measure photoelectrons in all possible directions. Since the measurement, the system, and the light are completely isotropic, there will be no net current in the system. However, if one introduces a spin detection axis, $\boldsymbol{\hat{s}}$, it breaks the symmetry and a net current along $\boldsymbol{\hat{s}}$ is symmetry allowed (see Fig. 1). The current conditioned on spin can only have a component along the spin detection axis. Since the current is a time-odd polar vector and spin is a time-odd axial vector, the projection of the current on the spin axis is a time-even pseudoscalar. Its pseudoscalar nature dictates that such current can only emerge in chiral molecules; as a time even quantity, it does not require breaking of the time-reversal symmetry of the Hamiltonian. Thus, the current is absent without spin conditioning and emerges as soon as the measurement is conditioned (or postselected) on spin.
To address the strength and nature of the these pseudoscalar correlations, we use our geometric formalism developed for spin-resolved one-photon photoionization of chiral molecules [13]. It allows us to identify and quantify the molecular pseudoscalrs underlying both time-even spin-current correlations as well as time-odd triple correlations between the spin of the photoelectron, the photoelectron current, and the photon spin.
@@
We show that these correlations are enabled by two complementary geometric mechanisms that arise solely from electric dipole interactions and the geometric properties of the photoionization dipoles in real space and spin space. The first mechanism is mediated by the momentum-resolved Bloch pseudovector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ defined below. It operates in fields of arbitrary polarization, including fully isotropic polarization. The second, time-odd, mechanism is mediated by the spin-resolved propensity field $\vec{\mathbb{B}}_{\boldsymbol{\vec{k}}}$ which is a natural extension of its spin-averaged precursor [25, 3] responsible for photoelectron circular dichroism [30, 27, 4, 21, 18] and is also defined below. This molecular property requires breaking time-reversal symmetry and can be activated by photon spin.
@@
To quantify our results we use synthetic chiral argon system employed in Ref. [13], which are constructed by combining excited-state orbitals:
@@
These sates $|\psi_{m,\mu}^{\pm}\rangle_{p}$ and $|\psi_{m,\mu}^{\pm}\rangle_{c}$ are inspired by analogous chiral hydrogenic states [25]. Unlike hydrogen, the multielectron core potential in argon breaks inversion symmetry: the synthetic chirality is stabilized by electron correlations.
@@
The photoelectron current conditioned on the spin detection axis $\boldsymbol{\hat{s}}^{L}$ can then be calculated as follows:
@@
wherein, $\int d\rho$, $\int d\Theta_{k}^{L}$, and $\int d\Theta_{s}^{L}$ denotes averaging over all molecular orientations, photoelectron momentum, and spin detection axis, respectively. Full details of the calculations are presented in the Supplementary Material.
[…]
Let us first consider randomly oriented chiral molecules under isotropic illumination by linearly polarized light:
$\boldsymbol{\vec{E}}^{L}=$ $E_{\omega}^{L}\left(\sin\theta_{p}\cos\varphi_{p}\boldsymbol{\hat{x}}^{L}+\sin\theta_{p}\sin\varphi_{p}\boldsymbol{\hat{y}}^{L}+\cos\theta_{p}\boldsymbol{\hat{z}}^{L}\right),$ (7)
where the angles $\theta_{p}$, $\varphi_{p}$ describe all possible polarization directions. Substituting Eq. (7) into (6) and averaging over all polarization directions $\int d\Theta_{p}$, the resulting photoelectron current is
@@
where,
$S_{0}=\int d\Theta_{k}\left(\left|\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},\frac{1}{2}}\right|^{2}+\left|\boldsymbol{\vec{D}}_{\boldsymbol{\vec{k}},-\frac{1}{2}}\right|^{2}\right),$ (9)
is proportional to the total ionization yield. Equation (8) shows that the current is proportional to the flux of the vector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}$ through the surface of the energy shell ($d\boldsymbol{\vec{\Theta}}_{k}^{M}=d\Theta_{k}^{M}\boldsymbol{\hat{k}}^{M}k^{2}$). Mathematically, $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ is a Bloch pseudovector:
@@
describing spin orientation in the degenerate two-level system formed by spin-up and spin-down continuum states of the photoelectron. It incorporates both populations and coherences within this system [13]. Here, $\tilde{\varrho}$ is the reduced density matrix which emerges after performing a trace over the degenerate ionization channels of the full density matrix, then averaging over all molecular orientations. Moreover, the Bloch vector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ defines a spin-texture on the energy shell for a fixed in space chiral molecule photoionized by isotropic illumination, see Fig. 2(a,b).
@@
and that the current is collinearly ‘locked’ to the direction of the spin-detection axis $\boldsymbol{\hat{s}}$: the direction of the current is collinear with spin detection axis. The strength of these correlations is defined by Eq. (8): the pseudoscalar describing spin-current correlations is a flux (or the net radial component) of the momentum-resolved Bloch vector. This expression shows the link between time-odd spin- molecular orientation correlations [13] driven by the projection of the net Bloch vector, $\boldsymbol{\vec{\mathcal{S}}}_{\text{net}}=\int d\Theta_{k}\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$, on to a molecular bond and its time-even spin-current correlations driven by the net radial component (or the flux of the momentum resolved Bloch vector through the energy shell). This result establishes momentum-resolved Bloch vector as a primary molecular property defining all pair-wise pseudoscalar correlations in photoionization of chiral molecules.
Figure 2c shows the strength of $\boldsymbol{\vec{j}}_{\text{iso}}^{L}$ for the chiral states $|\psi_{-1,\frac{1}{2}}^{\pm}\rangle_{c}$ and $|\psi_{1,\frac{1}{2}}^{\pm}\rangle_{p}$. The former shows that it can reach up to $3\%$ of the total signal while the latter shows that it is negligible. From Fig. 2(a) we see that the flux of $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ is mostly positive (negative) for the state $|\psi_{-1,\frac{1}{2}}^{+}\rangle_{c}$ ($|\psi_{-1,\frac{1}{2}}^{-}\rangle_{c}$) which is consistent with the sign of $\boldsymbol{\vec{j}}_{\text{iso}}^{L}$ in Fig. 2c. For the state $|\psi_{1,\frac{1}{2}}^{\pm}\rangle_{p}$, there is almost the same amount of positive and negative flux of the Bloch vector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ which results in a negligible $\boldsymbol{\vec{j}}_{\text{iso}}^{L}$.
The resulting photoelectron current can by enhanced by having a fixed polarization direction in the laboratory frame. Repeating the same process for linearly polarized,
@@
The additional vector $\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}$ is the directional bias introduced by the well-defined direction of the light polarization. Hence, the photoelectron current now arises as the flux of an effective Bloch pseudovector $\boldsymbol{\vec{\mathsf{S}}}_{\boldsymbol{\vec{k}}}=a\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}+b\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}$ through the surface of the energy shell. The current $\boldsymbol{\vec{j}}_{\text{lin}}^{L}$, Eq. (13), can be measured using two possible detection geometries: (i) collinear $\boldsymbol{\hat{s}}\parallel\boldsymbol{\hat{\epsilon}}$ and (ii) orthogonal $\boldsymbol{\hat{s}}\perp\boldsymbol{\hat{\epsilon}}$, see Fig. 2c. Enhancement or reduction of $\boldsymbol{\vec{j}}_{\text{lin}}^{L}$ with respect to $\boldsymbol{\vec{j}}^{L}_{\text{iso}}$ is then controlled by $\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}$:
@@
The comparison of the currents $\boldsymbol{\vec{j}}_{\text{iso}}^{L}$ and $\boldsymbol{\vec{j}}_{\text{lin}}^{L}$ is shown in Fig. 2c. It can be seen that the changes in the current are sensitive to both the chiral state and detection geometry.
Figure 3: Schematic of the resulting photoelectron currents for circularly polarized light in the laboratory frame. The solid and dashed blue vectors denote the opposite directions of spin for opposite enantiomers. (a,b) The current $\boldsymbol{\vec{j}}_{r}$ is enabled by the Bloch vector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ which results in a collinear locking of the photoelectron current and spin $\boldsymbol{\vec{j}}\parallel\boldsymbol{\hat{s}}$. (c,d) The currents $\boldsymbol{\vec{j}}_{\text{PECD}}$ and $\boldsymbol{\vec{j}}_{\tau}$ are enabled by the spin-resolved propensity field $\vec{\mathbb{B}}_{\boldsymbol{\vec{k}}}$. The PECD current is not spin-sensitive, while $\boldsymbol{\vec{j}}_{\tau}$ presents a triple locking of the photoelectron current, its spin with the photon spin in three mutually orthogonal directions.
[…]
@@
The term $\boldsymbol{\vec{j}}_{r}^{L}$ is enabled by an Bloch pseudovector $\boldsymbol{\vec{\mathsf{S}}}_{\boldsymbol{\vec{k}}}$ and has a similar form as that of Eq. (13):
@@
which is enantio-sensitive but not dichroic. The collinear locking between the photoelectron current and spin is still present and can be measured either via collinear or orthogonal detection geometry with respect to the photon spin direction $\boldsymbol{\hat{\Xi}}^{L}=(-i\boldsymbol{\vec{E}}^{L*}\times\boldsymbol{\vec{E}}^{L})/|\boldsymbol{\vec{E}}^{L}|^{2}=\xi\boldsymbol{\hat{z}}^{L}$, see Fig. 3(a,b). Moreover, the current $\boldsymbol{\vec{j}}_{r}^{L}$ is contained in the coefficients $B_{1}$ and $B_{2}$ introduced by Cherepkov [7, 8, 10]. Note that the same pseudoscalars (the fluxes of $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}^{M}$ and $\boldsymbol{\vec{\mathcal{S^{\prime}}}}_{\boldsymbol{\vec{k}}}^{M}$) contribute with different coefficient determined by the detection geometry.
@@
which is now a matrix in spin space. Explicitly, we have
@@
which yields the familiar PECD (photoelectron circular dichroism [27, 15, 35, 17, 21, 14, 18, 19, 22, 33, 9, 28, 34, 2]) current which is not spin sensitive, presenting a “background” in a spin-conditioned measurement. Equation (19) is equal to the PECD current we derived in [24], to the original expression derived by Ritchie [30], and to the coefficient $D$ introduced by Cherepkov [7, 8, 10].
@@
is proportional to the flux of the spin torque vector $\boldsymbol{\vec{\tau}}_{\boldsymbol{\vec{k}}^{M}}^{M}$ through the surface of the energy shell. This current is maximal when the photoelectron spin detection axis $\boldsymbol{\hat{s}}^{L}$ is orthogonal to both the direction of the current and photon spin $\boldsymbol{\hat{\Xi}}^{L}$. Thus, it is confined in the polarization plane and the respective spin-current correlations are transversal. The momentum-resolved spin-torque vector $\boldsymbol{\vec{\tau}}_{\boldsymbol{\vec{k}}^{M}}^{M}$ quantifies the intrinsic coupling between the photon and photoelectron spin mediated by the spin-resolved propensity field. $\boldsymbol{\vec{\tau}}_{\boldsymbol{\vec{k}}^{M}}^{M}$ defines the direction of the torque exerted on the photoelectron spin by the photon spin for every direction of the photoelectron momentum in the molecular frame. In this frame, the photon spin does not point along a single direction but instead forms a radially symmetric (hedgehog-like) vector field. This geometry enables access to the intrinsic photon–electron spin coupling without directional bias from the photon spin. $\boldsymbol{\vec{\tau}}_{\boldsymbol{\vec{k}}^{M}}^{M}$ exhibits skyrmionic texture (Fig. 4) which will be elucidated in our future work .
Equation (20) is also equivalent to the coefficient $C$ of Cherepkov [7, 8, 10]. The enantio-sensitive nature of $\boldsymbol{\vec{j}}_{\tau}^{L}$ arises from the molecular factor $\boldsymbol{\vec{\tau}}^{M}\cdot\boldsymbol{\vec{k}}^{M}$ (its net value, i.e. averaged over the orientation of $\boldsymbol{\vec{k}}^{M}$, see Eq.20) which is a pseudoscalar: opposite enantiomers produce opposite spin vortices, such that the spin polarization direction depends on the photoelectron momentum in a way reminiscent of Rashba effect in solids.